You have a choice of two bonds X and Y. Bond X is a $1,000 par value30-year 8% semiannual coupon government bond selling at $920. Bond Y is a $1,000 par value 20-year 10% semiannual coupon corporate bond selling at $1,057. You believe that the yield curve for government securities will be flat after 5 years at 7.5% compounded semiannually, while that for the corporate bond Y will be 8.5% compounded semiannually after 5 years. If you can re-invest coupons at 3.5% per half-year in the coming 5 years, what bond would you choose over an investment horizon of 5 years?
Answer #1

Expected holding period yield for Bond X=RATE(5,0,-920,FV(3.5%,5*2,-8%*1000/2)+PV(7.5%/2,25*2,-8%*1000/2,-1000))=10.64%

Expecte holding period yield for Bond Y=RATE(5,0,-1057,FV(3.5%,5*2,-10%*1000/2)+PV(8.5%/2,15*2,-10%*1000/2,-1000))=10.13%

Choose Bond X as it offers higher expected yield